OpenStudy (anonymous):

Write an equation for a line parallel to x+15=-5y passing through (0,4)

OpenStudy (mathstudent55):

Parallel lines have the same slope, First, you need to find the slope of the given line, then use that slope for the new line. To find the slope of the given line, solve x + 15 = -5y for y. Then put it in slope-intercept form, which is y = mx + b. m is the slope. Second, the equation of a line with slope m that passes through point (x1, y1) is y - y1 = m(x - x1) Try it, and ask for help if you need it.

OpenStudy (anonymous):

Thanks,I need help putting x=15=-5y into slope intercept form

OpenStudy (mathstudent55):

First, switch the sides, so you start with -5y on the left side.

OpenStudy (mathstudent55):

Next, the only thing happening to the y is that it is being multiplied by -5. How do you undo a multiplication?

OpenStudy (whpalmer4):

\[x+15 = -5y\]Divide both sides by -5 to give\[-\frac{x}{5} -3 = y\]Now switch the sides.

OpenStudy (anonymous):

Umm I think that by dividing?? So will that mean that I'm dividing 15 by -5y??

OpenStudy (mathstudent55):

Correct, divide both sides by -5

OpenStudy (anonymous):

So then what will happen next to the x??

OpenStudy (mathstudent55):

-5y = x + 15 Divide -5y by -5 Divide x by -5 Divide 15 by -5

OpenStudy (anonymous):

OK I got -5-3=y ..?..?..?

OpenStudy (anonymous):

wait wait! I checked it and it actually its suppose to be y=-1/5-3

OpenStudy (mathstudent55):

-5y/(-5) = x/(-5) + 15/(-5) y = -(1/5)x - 3

OpenStudy (mathstudent55):

Remember to include the x on the right side.

OpenStudy (anonymous):

YESSSSSSS!!!!!

OpenStudy (mathstudent55):

Since the original equation is now written in slope-intercept form, y = mx + b, you can read the slope directly. It's the number multiplying x, which is m. m = ?

OpenStudy (anonymous):

m=4 so then that will mean that the final answer will be y=-1/5x+4 xD

OpenStudy (mathstudent55):

Where did you get a 4 from y = -(1/5)x - 3?

OpenStudy (mathstudent55):

Remember, the slope-intercept form is y = mx + b, where m, the number multiplying x, is the slope.

OpenStudy (anonymous):

ohh ok then my bad

OpenStudy (mathstudent55):